# Evaluating the Utility of Conformal Prediction Sets for AI-Advised Image Labeling

### v1

@@ 3.2.1. Overview of Stimuli Generation
To generate stimuli, we sourced 80 task images from ILSVRC 2012, equally distributed between in-distribution and OOD (i.e., 40 in each). Each set of stimuli is further categorized by difficulty (between easy and hard) and set size (between small and large). This results in four distinct groups, each containing ten images, for both in-distribution and OOD. Image stimuli in each group were selected to exemplify their respective categories, using the model’s (described below) prediction confidence and the size of the derived prediction set.
@@ 3.2.2. Predictive Model for Image Classification
Given that underperforming models are rarely deployed in real-world scenarios, we deployed a classifier designed to achieve high prediction accuracy. We use a pre-trained Wide Residual Network (WRN) (Zagoruyko and Komodakis, 2016) on ILSVRC 2012. The WRN is a Residual Network (ResNet) variant (He et al., 2016), and its primary architectural distinction is the widening factor in the channel width of convolutional layers compared to a ResNet of the same depth. Specifically, we utilize a wide_resnet101_2 model with layers and a widening factor of , using PyTorch’s IMAGENET1K_V2 weights (Paszke et al., 2019). This model is well calibrated for in-distribution instances—achieving a Top- accuracy of and a Top- accuracy of over all in-distribution images, as shown in Table S1 of Appendix A. We use WRN to create Top- and Top- predictions and their corresponding softmax pseudo-probabilities to establish prediction displays for Top- conditions.
@@ 3.2.3. Calibration Set
To produce prediction sets, we randomly sample half of the 50,000 images from the ILSVRC 2012 validation set to create a conformal calibration set, which contains 25,000 in-distribution images that are unseen by the model. This calibration set is used to perform inductive conformalization by calibrating the model based on the conformity scores derived from all images of the conformal calibration set, which enables the creation of prediction sets for the remaining 25,000 images in the conformal hold-out set. Our prediction sets achieve size-stratified conditional coverage (), as shown in Table 1. It implies that, on average, there is a probability that an image’s true label class falls within the prediction set, conditioned on all set sizes.
| Size | Count | Coverage |
@@
We use OOD images to allow comparing the utility of different prediction displays in cases where the classifier is poorly calibrated, which lowers the accuracy of Top- predictions and increases the size of prediction sets. To create OOD images, we take a systematic approach to transform in-distribution images from our conformal hold-out set to OOD by applying covariate shifts (Quinonero-Candela et al., 2008) through synthetic image corruption (Michaelis et al., 2019). In supervised learning, where , the joint distribution over images and labels can be represented as , covariate shift refers to alterations to , the probability of the features , without altering the classifier .
After experimenting with 15 corruption types introduced by Michaelis et al. (2019), we opted to use five corruptions that reduced WRN’s Top- prediction accuracy the most. These corruptions are: (1) defocus blur, (2) snow, (3) frost, (4) zoom blur, and (5) glass blur, which are bolded in Table S1 of Appendix A. Given that the human visual system can be resilient to certain minor changes (Michaelis et al., 2019; Hendrycks and Dietterich, 2019), we corrupt all images in the conformal hold-out set at the maximum severity, as defined by Michaelis et al. (2019).
Table S1 shows that our prediction sets achieve a coverage rate for in-distribution images (Vovk et al., 2005; Shafer and Vovk, 2008). However, the coverage guarantee becomes invalid for OOD samples (Angelopoulos et al., 2020) as images used during calibration and testing are no longer exchangeable (i.i.d.).
@@
To compare the effects of the different prediction displays on images of varying difficulty, we categorize images by difficulty using cross-entropy loss, which calculates how well the predicted probability distribution aligns with the ground truth distribution. Formally in Equation 1, let be an input image; denotes the true probability distribution of the image’s label in one-hot encoding, while is our classifier’s predicted distribution.
Figure S1 in Appendix A displays the distribution of cross-entropy loss of all in-distribution images, with the median presented as an orange dotted line. Due to WRN’s high predictive power, this distribution is highly right-skewed, which creates a problem: if we use this median threshold to categorize difficulty, we would include many images in the easy group that the model is both confident and accurate in predictions, so the derived prediction sets embody little uncertainty, synonymous to Top- prediction. We thus compute a new median threshold () after excluding images with a set size of one such that the difficulty threshold, shown as a red solid line, is defined based on images with non-deterministic sets. Images with cross-entropy loss falling below or on this threshold are classified as easy, while those above the threshold are deemed hard.
Because the cross-entropy loss derived from the predictive distribution of OOD images can be misleading as a signal of task difficulty due to model misspecification (Guo et al., 2017; Torralba and Efros, 2011), we use the red median threshold from the in-distribution data to categorize the difficulty of OOD images.
@@
Images within each difficulty group of the conformal hold-out set can induce different prediction set sizes. For example, although images classified as easy are more likely to have a small set size, there are easy instances that require a large set size to achieve the desired coverage. Because we wish to understand how set size can impact the efficacy of prediction sets, we categorize stimuli within each difficulty group by size. After grouping images in each conformal hold-out set by easy and hard, we distinguish between images based on relative set size as produced by RAPS, categorizing them as either small or large, based on the median of the group. However, as shown in Table 2, the predictive power of WRN consistently demonstrates high confidence in its predictions for many easy images with a set size of one, which significantly lowers the median size threshold. To ensure more variance in set size within the easy groups, we compute the median threshold after excluding images with a set size of one (i.e., these are directly assigned to the small group). Images with a set size smaller than or equal to the size median are classified as small; if the opposite is true, they are classified as large. The specific thresholds used for each difficulty category within the conformal hold-out set are outlined in Table 2, with the bold figures representing the applied thresholds.
@@
Upon categorizing all images in each conformal hold-out set (i.e., one in-distribution and five OOD, totaling 6 25,000 images) by difficulty (easy and hard) and size (small and large), our objective is to select a set of task images from which to sample image assignments for participants, so as to avoid results that are overfit to a very small set of stimuli. Specifically, we target 80 task images from the 1,200 candidate images. Our goals are to ensure: (1) the candidate images are representative of each group defined by in-distribution or OOD, difficulty, and size (), and (2) task images avoid the known data anomalies in ILSVRC 2012 (Ekambaram et al., 2017; Northcutt et al., 2021).
Example stimuli used for the experiment.
@@
Sampling images from groups, such as in-distribution, easy (difficulty), and small (set size), is challenging due to the prevalence of images with a set size of one, for which prediction sets become nearly identical to the Top- prediction. To avoid oversampling sets of size one, (1) we temporarily set aside images with a set size of one, (2) we take images with sets larger than one whose set size falls near the median size in that group (i.e., and percentiles), and (3) perform a post hoc sampling correction to include images with a set size of one in proportion to their original presence in each group, as shown in Table S2 of Appendix A.
@@
Images from ILSVRC 2012 are known to be noisy, with numerous data anomalies. To create the set of 80 task images, we thereby sample 50 candidate task images from each representative subset defined by in-distribution or OOD, difficulty, and size () without replacement. We manually inspect the task images, starting from images with higher cross-entropy loss (hard), and omit those with (1) an obviously wrong ground truth label, (2) a text overlay with the correct label, (3) no obvious focal object or many surrounding objects, (4) unusual dimensions, and (5) small multiples. Based on these rules, we selected 10 images from each of the 4 groups generated by the combination of that are in-distribution, and 2 images from each of the 4 groups with corruptions. This process generates 80 task images in total, balanced by in-distribution or OOD, difficulty, and set size.
Table 3 presents the Top- accuracy, coverage, and mean set size grouped by our experimental factors with example stimuli shown in Figure 2. Our stimuli achieve the desired behavior, as the average prediction accuracy and set size vary according to the difficulty and set size manipulated. Notably, despite Top- prediction accuracy being lower for OOD stimuli, prediction sets for OOD stimuli maintain a high coverage rate. This is because the adaptiveness of the conformal prediction set (Romano et al., 2020; Angelopoulos et al., 2020) makes them more resilient to random perturbations from image corruptions. The prediction set can safeguard the coverage for certain instances by significantly increasing the set size, as shown by the bolded size summary in Table 3.
| Shift | Difficulty | Size | Count | acc@1 | acc@10 | Coverage | Avg. Set Size |
| in | easy | small | 10 | 1 | 1 | 1 | 2.6 |
| large | 10 | 1 | 1 | 1 | 19.5 | ||
| hard | small | 10 | 0.4 | 1 | 0.9 | 8.0 | |
| large | 10 | 0.5 | 0.9 | 0.9 | 51.1 | ||
| out | easy | small | 10 | 1 | 1 | 1 | 5 |
| large | 10 | 1 | 1 | 1 | 28.1 | ||
| hard | small | 10 | 0 | 0.6 | 0.9 | 30.1 | |
| large | 10 | 0.1 | 0.2 | 0.9 | 90.8 |
@@ 3.2.4. Reward and Bonus
Participants earn a base compensation of for completing the study and get a bonus for each correct label. With 16 trials, the maximum bonus is , bringing the total potential earnings to . We determined the task reward to achieve local minimum wage based on the assumption that most participants would take longer than the average time we observed in a small pilot of the experiment.
@@ 3.3. Experimental Interface
Our task interface222Study interface: https://conformal-prediction.com, of which screenshots are presented in Figure 3, uses a two-column design, presenting task images on the left and predictions on the right, with a response field directly below the task image. In the baseline condition without predictions, the task image and response field are centered. Participants can see real-time earnings in the top-right corner of the interface.
To ensure that participants understood the meaning of each label, we selected label-representative images from the conformal calibration set. To find suitable label-representatives, we manually examined all images by labels and prioritized selection for those with small cross-entropy loss (i.e., easy), following the same rules used when selecting task images. The label-representative images we used are included in the Supplemental Material.
@@ 3.3.1. Presenting Predictions
For the three treatment conditions that feature predictions, the Top- condition presents participants with the classifier’s Top-1 prediction and its softmax score on top of a representative image of the predicted label. Top- predictions with softmax scores are ordered in a table, while the prediction set is presented in an -by- matrix sorted column-wise by the classifier’s prediction confidence. Participants could hover their mouse cursor over any predicted label to view a label-representative image.
Interface screenshots by conditions.
@@ 3.3.2. Label Space Search
To label task images, participants need to be aware of the label space. Because ILSVRC 2012 has a large label space of 1,000, which is cumbersome to present in a dropdown or checkbox format, we organized labels using a version of the WordNet hierarchy, which we truncated slightly to avoid label overlap. For example, automobile and truck are both parent nodes for minivan, and since truck is also a type of automobile, we keep minivan in the automobile category. We include this tree in the Supplemental Material.
If participants disagree with the predictions or when no prediction is available (baseline), our interface enables them to perform three types of search. As illustrated in Figure 4, they are: (A) dropdown search, (B) keyword search, and (C) bottom-up search. In dropdown search, a dropdown menu appears as participants begin typing in the response field, suggesting leaf labels that match their input. Alternatively, in keyword search, participants can type and then click a search button to search the hierarchy for specific keywords. Doing so opens an overlay window showing the label-space hierarchy network, where leaf nodes (denoting valid labels) appear in a different color from the parent nodes. Clicking on a leaf node populates the label in the response field. To ease navigation, the hierarchy network is collapsed by default except for the parent nodes (i.e., categories) and leaf nodes (i.e., labels) containing the search keywords with edges connecting them highlighted in orange. All other parent nodes can also be clicked to expand.
Participants can also perform a bottom-up search starting from the prediction display. Clicking on a predicted label opens the search overlay. Instead of showing the whole hierarchy network, the overlay presents the path from the root node to the leaf node corresponding to the predicted label the participant clicked. Participants can explore adjacent labels under the same parent category or explore other parent nodes on the path upward toward the root node.
When exploring the label space hierarchy, participants can hover over any leaf node to display its label-representative image. We illustrate these search actions through video demonstrations included in the Supplemental Material.
Interface screenshots by conditions.
@@ 3.4. Experimental Procedure
Participants are directed to our study interface from Prolific. On the welcome page, they receive a brief description of the study’s purpose and the estimated time required for completion. Participants are instructed to complete the study in a single session using Google Chrome on a large-screen device. If participants agree to the terms by clicking the “Start” button, they are randomly assigned to a prediction display condition with task images sampled from each corresponding group, shuffled in a randomized order.
Participants then review two instruction pages. The first introduces the interface layout and details the bonus mechanism. The second contains several short videos, and demonstrates how to explore predictions (except in the baseline condition) and navigate the label space. Participants must watch all provided instructional videos before proceeding to an example page where they can practice using the interface. Then, they complete 16 rounds of tasks. Upon finishing the 16 trials, participants are asked to state their willingness-to-pay to have access to the same style of prediction display they used in the experiment on a scale from to USD (i.e., the range of possible earnings from the experiment) if they were invited to undertake another 16 rounds of tasks with new images generated similarly to those they had experienced, for further reward and potential bonus. On the same page, participants could optionally describe their strategy to identify the correct label for each image, with or without aid from predictions.

### v6

@@ 3.2.1. Overview of Stimuli Generation
To generate stimuli, we sourced 80 task images from ILSVRC 2012, evenly divided between in-distribution and OOD (i.e., 40 of each). For both in-distribution and OOD images, we further categorized each set of stimuli by difficulty (between easy and hard) and by set size (between smaller and larger). Note that because the range of observed set sizes differed depending on the difficulty of the instance, set sizes were defined relatively rather than absolutely. This categorization resulted in four distinct groups, each comprising ten images, for both in-distribution and OOD categories. Image stimuli within each group were selected to exemplify their respective categories, determined by the model’s (described below) prediction confidence and the size of the prediction set derived.
@@ 3.2.2. Predictive Model for Image Classification
Given that underperforming models are rarely deployed in real-world scenarios, we chose a classifier designed to achieve high prediction accuracy. We used a pre-trained Wide Residual Network (WRN) (Zagoruyko and Komodakis, 2016) on ILSVRC 2012. The WRN is a Residual Network (ResNet) variant (He et al., 2016), and its primary architectural distinction is the widening factor in the channel width of the convolutional layers compared to a ResNet of the same depth. Specifically, we utilized a wide_resnet101_2 model with layers and a widening factor of , using PyTorch’s IMAGENET1K_V2 weights (Paszke et al., 2019). This model was well-calibrated for in-distribution instances, achieving a Top- accuracy of and a Top- accuracy of over all in-distribution images, as detailed in Table 6 of Appendix A. We used the WRN to create Top- and Top- predictions along with their corresponding softmax pseudo-probabilities to establish prediction displays for Top- conditions.
@@ 3.2.3. Calibration Set
To produce prediction sets, we randomly sampled half of the 50,000 images from the ILSVRC 2012 validation set to create a conformal calibration set, which contained 25,000 in-distribution images that were unseen by the model. This calibration set was used to perform inductive conformalization by calibrating the model based on the conformity scores derived from all images in the conformal calibration set. This process enabled the creation of prediction sets for the remaining 25,000 images in the conformal hold-out set. Our prediction sets achieved size-stratified conditional coverage (), as shown in Table 1. This implies that, on average, there is a probability that an image’s true label class falls within the prediction set, conditioned on all set sizes.
| Size | ||
| Count | Coverage | |
@@
To compare the utility of different prediction displays when the classifier is poorly calibrated, we utilized OOD images, which would lower the accuracy of Top- predictions and increase the size of prediction sets. To create OOD images, we systematically transformed in-distribution images from our conformal hold-out set into OOD by applying covariate shifts (Quinonero-Candela et al., 2008) through synthetic image corruption, as described by Michaelis et al. (2019). Covariate shift, in the context of supervised learning where the joint distribution over images and labels is , refers to changes in , the feature probability, without affecting the classifier .
After evaluating 15 types of corruption, we opted to use five corruptions that reduced WRN’s Top- prediction accuracy the most. These corruptions were: (1) defocus blur, (2) snow, (3) frost, (4) zoom blur, and (5) glass blur, which are bolded in Table 6 of Appendix A. Given that the human visual system can be resilient to certain minor changes (Michaelis et al., 2019; Hendrycks and Dietterich, 2019), we applied these corruptions to maximum severity for all images in the conformal hold-out set, as defined by Michaelis et al. (2019).
Table 6 of Appendix A indicates that our prediction sets achieve a coverage rate for in-distribution images. However, the coverage guarantee for the OOD samples was compromised, as the images used during calibration and testing are no longer exchangeable (i.i.d.), according to Angelopoulos et al. (2020).
@@
To compare the effects of different prediction displays on images of varying difficulty, we categorized images by difficulty using the cross-entropy loss, which calculates how well the predicted probability distribution aligns with the ground truth distribution. Formally in Equation 1, let be an input image; denotes the true probability distribution of the image’s label in one-hot encoding, while is our classifier’s predicted distribution.
Figure 9 of Appendix A displays the distribution of cross-entropy loss of all in-distribution image stimuli, with the median presented as an orange dotted line. Due to WRN’s high predictive power, this distribution was highly right-skewed, which created a problem: if we used this median threshold to categorize difficulty, we would include many images in the easy group that the model was both confident and accurate in predictions, so the derived sets embodied little uncertainty, synonymous with Top- prediction. We thus computed a new median threshold () after excluding images with a set size of one such that the difficulty threshold, shown as a red solid line, was defined based on images with non-deterministic sets. Images with cross-entropy loss falling below or on this threshold were classified as easy, while those above were deemed hard.
Because the cross-entropy loss derived from the predictive distribution of OOD images can be misleading as a signal of task difficulty due to model misspecification (Guo et al., 2017; Torralba and Efros, 2011), we used the red median threshold from the in-distribution data to categorize the difficulty of OOD images.
@@
Images within each difficulty group of the conformal hold-out set induce varied prediction set sizes. For instance, while images classified as easy typically had a smaller set size, some easy instances required a larger set size to achieve the desired coverage. To explore how set size influences the efficacy of prediction sets, we categorized the stimuli within each difficulty group by their set size. After grouping images in each conformal hold-out set by easy and hard, we further distinguished them based on relative set size, as determined by RAPS, into smaller or larger categories, based on the median of the group. However, as shown in Table 2, the predictive power of WRN consistently demonstrated high confidence in its predictions for many easy images with a set size of one, which significantly lowered the median size threshold. To introduce more variance in set size within the easy categories, we calculated the median threshold after excluding images with a set size of one (i.e., these were directly assigned to the smaller group). Images with a set size smaller than or equal to this adjusted size median were classified as smaller; if the opposite is true, they were classified as larger. The specific thresholds used for each difficulty category within the conformal hold-out set are outlined in Table 2, with the bold figures representing the applied thresholds.
@@
Upon categorizing all images in each conformal hold-out set (i.e., one in-distribution and five OOD, totaling 6 25,000 images) by difficulty (easy and hard) and size (smaller and larger), our objective was to select a set of task images from which to sample image assignments for participants, so as to avoid results that overfit to a very small set of stimuli. Specifically, we targeted 80 task images from the 1,200 candidate images. Our goals were to ensure that: (1) the candidate images are representative of each group defined by in-distribution or OOD, difficulty, and size (), and (2) task images avoid the known data anomalies in ILSVRC 2012 (Ekambaram et al., 2017; Northcutt et al., 2021).
| Shift | Difficulty | Size | Count | Coverage | Avg. Set Size | ||
| Top-1 | Top-10 | RAPS | |||||
| in | easy | smaller | 10 | 1 | 1 | 1 | 2.6 |
| larger | 10 | 1 | 1 | 1 | 19.5 | ||
| hard | smaller | 10 | 0.4 | 1 | 0.9 | 8.0 | |
| larger | 10 | 0.5 | 0.9 | 0.9 | 51.5 | ||
| out | easy | smaller | 10 | 1 | 1 | 1 | 5 |
| larger | 10 | 1 | 1 | 1 | 28.1 | ||
| hard | smaller | 10 | 0 | 0.6 | 0.9 | 30.1 | |
| larger | 10 | 0.1 | 0.2 | 0.9 | 90.8 | ||
This figure displays a grid showcasing example stimuli participants viewed for labeling tasks. Rows differentiate in-distribution and OOD stimuli, while columns further categorize stimuli examples by difficulty and set size.
@@
Sampling images from groups, such as in-distribution, easy (difficulty), and smaller (set size), was challenging due to the prevalence of images with a set size of one for which prediction sets become nearly identical to the Top- prediction. To avoid oversampling sets of size one, (1) we temporarily set aside images with a set size of one, (2) we took images with sets larger than one whose set size falls near the median size in that group (i.e., th and th percentiles), and (3) we performed a post-hoc sampling correction to include images with a set size of one in proportion to their original presence in each group, as shown in Table 8 of Appendix A.
@@
Images from ILSVRC 2012 are known to be noisy, with numerous data anomalies. To create the set of 80 task images, we sampled 50 candidate task images from each representative subset defined by in-distribution or OOD, difficulty, and size () without replacement. We manually inspected the task images, starting from images with higher cross-entropy loss (hard), and omitted those with (1) an obviously wrong ground truth label, (2) a text overlay with the correct label, (3) no apparent focal object or many surrounding objects, (4) unusual dimensions, and (5) small multiples. Based on these rules, we selected 10 images from each of the 4 groups from the combination of that were in-distribution, and 2 images from each of the 4 groups with corruption. This process generated 80 task images in total, balanced by in-distribution or OOD, difficulty, and set size.
This figure contains screenshots of the interface used by participants to complete the study, organized by conditions: baseline, Top-1, Top-10, and RAPS. Each screenshot of the interface illustrates the layout, adapted to accommodate different prediction displays with varying uncertainty quantification.
Table 3 presents the Top- accuracy, coverage, and average set size, grouped by our experimental factors, with example stimuli shown in Figure 2. Our stimuli exhibited the desired behavior, with the average prediction accuracy and set size varying in response to the manipulated difficulty and set size. While the stimuli we selected to achieve balance over our experimental manipulations resulted in close conformal coverage for OOD instances to the in-distribution case, anecdotally we found that OOD coverage tended to remain high () over varying stimuli sets under our data generating model.
@@ 3.2.4. Reward and Bonus
Participants earned a base compensation of for completing the study and received a bonus for each correct label. With 16 trials, the maximum bonus was , bringing the total potential earnings to . We determined the task reward to achieve the local minimum wage based on the assumption that most participants would take longer than the average time we observed in a small pilot of the experiment.
@@ 3.3. Experimental Interface
Our task interface, screenshots of which are presented in Figure 3, featured a two-column design: task images were on the left, and predictions were on the right, with a response field directly below the task image. In the baseline condition without predictions, the task image and the response field were centered. Participants could view their real-time earnings in the top-right corner of the interface.
To ensure that participants understood the meaning of each label, we selected label-representative images from the conformal calibration set. To find suitable label-representatives, we manually examined all images by labels and prioritized selection for those with small cross-entropy loss (i.e., easy), applying the same rules used when selecting task images. The label-representative images we used are included in the Supplemental Material.
@@ 3.3.1. Presenting Predictions
For the three treatment conditions that featured predictions, the Top- condition presented the classifier’s Top-1 prediction and its softmax score on top of a representative image of the predicted label. Top- predictions with softmax scores were ordered in a table, while the prediction set was presented in an -by- matrix sorted column-wise by the classifier’s prediction confidence. Participants could hover their mouse cursor over any predicted label to view a label-representative image.
This figure contains screenshots to demonstrate the search features of the interface provided to participants to assist with their labeling tasks. When performing a dropdown search, a dropdown menu will appear below the input field, which shows potential labels that match what the participants typed. Keyword search enables participants to search for labels that match their typed keywords by clicking a “Search” button. Bottom-up search allows participants to explore the label space by clicking on a predicted label. Additionally, participants can hover over any leaf node in the hierarchy network to access label-representative images.
@@ 3.3.2. Label Space Search
To label task images, participants need to be aware of the label space. Because ILSVRC 2012 has a large label space of 1,000, cumbersome to present in a dropdown or checkbox format, we organized labels using a version of the WordNet hierarchy, which we truncated slightly to avoid label overlap. For example, “automobile” and “truck” are both parent nodes for “minivan” and since “truck” is also a type of “automobile”, we retained “minivan” in the “automobile” category. We include this tree in the Supplemental Material.
If participants disagreed with the predictions or when no prediction was available (baseline), our interface enabled them to perform three types of search. As illustrated in Figure 4, they were (A) dropdown search, (B) keyword search, and (C) bottom-up search. In the dropdown search, a dropdown menu would appear as participants began typing into the response field, suggesting leaf labels that matched their input. Alternatively, in the keyword search, participants could type and then click a search button to search the hierarchy for specific keywords. Doing so would open an overlay window showing the label-space hierarchy network, where leaf nodes (i.e., denoting valid labels) would appear in a different color from the parent nodes. Clicking on a leaf node would populate the label in the response field. To ease navigation, the hierarchy network was collapsed by default except for the parent nodes (i.e., categories) and leaf nodes (i.e., labels) containing the search keywords with edges connecting them highlighted in orange. All other parent nodes could also be clicked to expand.
Participants could also perform a bottom-up search starting from the prediction display. Clicking on a predicted label would open the search overlay. Instead of showing the entire hierarchy network, the overlay presented the path from the root node to the leaf node corresponding to the predicted label that the participant clicked on. Participants could explore adjacent labels under the same parent category or explore other parent nodes on the path toward the root.
When exploring the label space hierarchy, participants could hover over any leaf node to display its label-representative image. We illustrate these search actions through video demonstrations included in the Supplemental Material.
@@ 3.4. Experimental Procedure
Participants were directed to our study interface from Prolific. On the welcome page, they received a brief description of the study’s purpose and the estimated time required for completion. Participants were instructed to complete the study in a single session using Google Chrome on a large-screen device. If participants agreed to the terms by clicking the “Start” button, they would be randomly assigned to a prediction display condition with task images sampled from each corresponding group, shuffled in a randomized order.
Participants then reviewed two instruction pages. The first introduced the interface layout and detailed the bonus mechanism. The second contained several short videos demonstrating how to explore predictions (except in the baseline condition) and navigate the label space. Participants had to watch all provided instructional videos before proceeding to an example page where they could practice using the interface. Then, they completed 16 rounds of tasks. Upon finishing the 16 trials, participants were asked to state their willingness-to-pay to have access to the same style of prediction display they used in the experiment on a scale from to USD (i.e., the range of possible earnings from the experiment) if they were invited to undertake another 16 rounds of tasks with new images generated similarly to those they had experienced, for further reward and potential bonus. On the same page, participants could optionally describe their strategy to identify the correct label for each image, with or without the aid of predictions.